# supplemental_arxiv

### v1

@@ 3 Quantile Gaussian process model
The covariance function $\kappa_{\theta}(\cdot,\cdot)$ initially appears as if it will be difficult to evaluate and indeed can be, depending on the functional forms of $f_{\theta}$ and $Q_{\theta}$. But where these functions belong to a location-scale family, the covariance function can be greatly simplified. Noting that the covariance is a function of $f_{\theta}(Q_{\theta}(p))=q_{\theta}(p)$, the QDF, if $\theta=(\mu,\sigma)$ where $\mu$ is a location parameter and $\sigma$ is a scale parameter, then $q_{\theta}(p)=\sigma q(p)$ where $q(p)$ is the QDF of a standard distribution having location 0 and shape 1. Staudte (2017) calls this the location invariant and scale equivariant property. For location-scale distribution families, this greatly simplifies the form of $\kappa_{\theta}(\cdot,\cdot)$ so that it becomes
@@ 3.1 Normal QGP
$\Psi_{ij}=\frac{2\pi(p\wedge p^{\prime}-pp^{\prime})}{\text{exp}\{-\frac{1}{2}[\Phi^{-1}(p)^{2}+\Phi^{-1}(p^{\prime})^{2}]\}}.$
@@
In this model, one simplifying assumption may be to treat $\sigma$ and $\gamma^{2}$ separately and proceed to fit a standard linear regression model. All frequentist and Bayesian results of the linear regression model apply, including the existence of conditionally conjugate prior distributions. The positive constraint on $\sigma$ can also be dealt with without adding much complication (see Gelman et al. (2013) pgs. 377-378). If, for a certain problem it is important to make inference on the unknown $n$, or $\sigma$ and $\gamma^{2}$ cannot treated as separate, then one may estimate the two parameters by assigning appropriate prior distributions to $\sigma$ and $n$ and reverting back to (10). In section 4, we analyze a normal QGP in a simulation study where we assign independent prior distributions to $\sigma$ and $n$, and $\sigma$ is treated as both a regression coefficient and as part of the variance.
@@ 3.2 Finite normal mixture QGP
In section 4, we analyze PIT transformed QGP model (12) where $F_{\theta}$ is set as a normal mixture distribution with $C=4$ component normal distributions. Fitting this model proved faster and produced better results than trying to fit the QGP model (9) and evaluating $Q_{\theta}$ the QF of a normal mixture. In fact we found that in cases where $F_{\theta}$ was a simpler function, such as a normal or exponential distribution, fitting model (12) was at least as fast and the results were at least as good as fitting (9). Thus when possible, we elected to fit (12).
@@ 3.3 Competing quantile matching methods
The SPL method was used by Gerding et al. (2023) and Shandross et al. (2024) in order to estimate a CDF function given quantile forecasts from disease outbreak forecast hubs. SPL is an interpolation where monotonic cubic splines are fit to pass through each given quantile and predict a function for all values between given quantiles and beyond the extreme values. An R package used for fitting SPL and which we use in section 4 is distfromq Ray and Gerding (2024).
The KDE method treats given quantiles as if they constitute a random sample from a distribution and applies kernel smoothing to estimate a density function. KDE requires selecting a kernel function, often a Gaussian kernel is chosen, and applying that function to each draw of a sample. Gyamerah et al. (2020) applied KDE smoothing with an Epanechnikov kernel to quantiles estimated via three different machine learning methods used for predicting crop yield. They then combined the estimated densities into an ensemble prediction. He et al. (2016) also use KDE to estimate density forecasts from quantiles to produce energy forecasts. We use the evmix package for performing QM using KDE Hu and Scarrott (2018).
One of the more common methods for QM is to select a CDF or QF function and then select model parameters which give a least squares fit to the estimated quantiles as done by Li et al. (2019). The R package rriskDistributions performs this least squares fit for some standard distributions Belgorodski et al. (2017). Independent random error may be included to the best fit QF allowing for estimation of the variability of the quantiles Nirwan and Bertschinger (2020). In section 4, we analyze model (13) as a proxy for the IND model to compare with other methods.
@@
The final method we include is the ORD model introduced by Nirwan and Bertschinger (2020). This model relies on the definition of sample quantiles being order statistics. In the ORD model, the joint distribution of a set of order statistics is the model likelihood. The likelihood of a set of order statistics from a continuous distribution is a function of both the CDF and PDF of a continuous distribution. ORD is the method most similar to QGP and it provides exact inference of quantile uncertainty whereas the QGP provides asymptotic inference of uncertainty.
@@ 3.4 Distance measures
A number of metrics for measuring the distance between two continuous univariate distributions exist. Those we consider in this paper are the Wasserstein distance (WD) and a slight modification of it, the total variation (TV) or the statistical difference, and the Kullback-Leibler divergence (KLD). A brief overview of these and other metrics, some of their statistical uses and properties, and relationships between metrics is found in Gibbs and Su (2002). These metrics were used in section 4 to assess the how well different QM methods approximate a true distribution.
The WD is used to measure the distance between two univariate random variables by their CDFs or QFs and has many applications in mathematics, optimization, and statistics, see Panaretos and Zemel (2019) for a review of the WD and its use in statistics. The $p$-WD is defined for two continuous random variables $X$ and $Y$ with respective CDFs $F_{X}$ and $F_{Y}$. In terms of the two CDFs the $p$-WD is defined in (14), and the $p$-WD in terms of the corresponding QFs is defined in (15).
[…]
The TV, defined in (17), is a distance between distributions measured in terms of the PDFs $f$ and $g$. The TV takes values between 0 and 1, and it is closely related to the distance used in Sgouropoulos et al. (2015) to measure the goodness of QM.
[…]
@@ 4 Quantile matching methods comparison on simulated data
In this section we present simulation studies assessing the QM of the QGP model and compare results with those of the SPL, KDE, IND, and ORD matching methods. The QGP, IND, and ORD models were each fit via Hamiltonian Monte Carlo (HMC) sampling using the Stan software and the cmdstanr package developed and maintained by the Stan Development Team (2024) Gabry et al. (2024). Simulation studies are done to assess parameter estimation and inference as well as QM estimation of a distribution.
4.1 Parameter estimation and quantile matching for known distribution
families
@@ 4.1.1 Normal parameter estimation
[…]
The top of figure 2 shows the 90% coverage of the posterior distribution credible intervals as $n$ increases for $\mu$, $\sigma$ and $n$ for five models. QGP-n and ORD-n models where $n$ is known are included along with QGP, ORD, and IND. The 90% credible intervals of the parameters were calculated by computing the $0.05^{th}$ and $0.95^{th}$ quantiles from the posterior distribution samples. The coverage percentage is calculated as the percentage of the 500 replicates which the true parameter value was within the 90% credible interval. As expected, the nominal coverage for the QGP and ORD models is better than that of the IND model, and for the two models where $n$ is known the coverage is better than where it is unknown, particularly for lower values of $n$ and $K$. The coverage for the ORD models appears to be slightly better than for the QGP models though not by much, and the difference decrases as $n$ increases. The superior performance of the ORD model is possibly because it provides exact inference whereas the QGP model provides asymptotic inference.
@@
To analyze how well QM methods produced predictive distributions which approximate the true distribution $N(4,3.5^{2})$, we measured the distance between the estimated predictive distributions of QGP, ORD, and IND fits to the true distribution. We also measured the distances of the SPL and KDE QM fits to the true distribution. The distances were measured via the UWD1, TV, and KLD metrics. The UWD1 for each of QGP, ORD, QGP-n, ORD-n, and IND were calculated as follows. For $\{\theta_{m}\}^{M}$ being $M$ draws from the posterior distribution of $\theta$ where $\theta_{m}=(\mu_{m},\sigma_{m})$, the QM posterior predictive distribution was simulated by repeatedly sampling a $\theta_{m}^{*}$, then sampling $X_{m}$ from a distribution with CDF $F_{\theta_{m}^{*}}$. Repeating this for $M=50,000$ times gives a QM posterior predictive sample. For the CDF $F_{\theta}$ where $\theta$ is the true parameter, $\xi_{m}=F_{\theta}(X_{m})$ was calculated and the empirical CDF $\hat{F}_{\xi}$ was calculated from the sample $\{\xi_{m}\}^{M}$. $F_{\xi}$ in (16) was then replaced by $\hat{F}_{\xi}$ to calculate UWD1. For SPL and KDE $F_{\xi}(x)$ was replaced by the respective estimated CDFs. The integrate function in R was used for calculating the UWD1 in (16).
@@
where $Y_{s}\overset{iid}{\sim}F_{\theta}$. For both SPL and KDE, $f_{\bar{\theta}_{M}}$ is the estimated PDF.
@@
This study shows the ability of the QGP QM model to estimate and perform inference on model parameters where the true distribution family is known. It also shows the QGP’s ability to produce a posterior predictive distribution which closely matches a true distribution, according to several metrics, relative to other QM methods. Between parameter inference and predicting the true distribution, QGP is superior to all methods we compared it with except for the ORD model which has similar results to the QGP. Additional plots for this study and for a similar simulation study where the data was simulated from an exponential family distribution are in the supplementary material. The results of the exponential family study are similar to those of the normal distribution family study.
@@
The analysis above show the usefulness in using the QGP model both in parameter estimation and in QM when given sample quantiles. In the normal case, the QGP showed improved parameter inference over the IND model and allows for making inference on the sample size $n$. Compared to the SPL and KDE QM methods, the QGP fits tend to be closer in distance to the true distribution. The QGP model, however, performs about the same as the ORD model in terms of inference and QM. The section below outlines a situation where the QGP model may be a better option for QM than the ORD model.
@@ 4.1.2 Quantile defined distributions
The GLD is reviewed in section 2.1.4, and a simulation carried out similarly to the previous studies was performed with GLD as the true distribution. However, QM was only done by the QGP and ORD models. Because of the relative difficulty of evaluating the CDF of the GLD and the ease of evaluating the QF, the CDF transformed QGP model in (9) instead of (12) was used. For the ORD model, evaluating the CDF was required for modeling. Evaluating the CDF of the GLD requires using an algebraic solver, and it took longer to code the ORD model in Stan due to a lack of readily available software which we were aware of. Once working, however, both QGP and ORD models fit the estimated quantiles well, however, the ORD usually required more time to fit than the QGP. Figure 10 in the supplementary material shows boxplots for the 500 replicates for different values of $K$ quantiles. In all cases except for $K=50$ the QGP tends to fit much faster than ORD. Note that the $y$-axis is on the logarithmic scale. Even at $K=50$ some outliers for the ORD model required hundreds of minutes to fit whereas the longest required time for fitting a QGP model was a few dozen minutes.
@@ 4.2 Matching unknown distributions
With the true distribution family being unknown, we made the modeling decision to set $F$ in the QGP model of (12) to be a mixture of normals distribution taking the form of (11). The number of component distributions was chosen to be $C=4$ to provide enough flexibility to approximate the different shapes of quantile functions. This decision was based on the common claim that any continuous distribution may be approximated by a finite mixture of normal distributions Peel and MacLahlan (2000); Nguyen and McLachlan (2019); Nguyen et al. (2020). This claim may be optimistic, but we found that a four component normal mixture distribution can approximate the non-normally shaped distributions reasonably well. However, any function which meets the technical definition of a CDF would be appropriate to use for $F$, see Gasthaus et al. (2019) for an example.
@@
We analyzed QM methods for quantiles from three different distributions; the extreme value (EV) distribution with location 0 and scale 1 $EV(0,1)$, Laplace (La) with location 0 and scale 1 $La(0,1)$, and a two component normal mixture (MIX) $wN(-1,0.9)+(1-w)N(1.2,0.6)$ where $w=0.35$. From each the three distributions, 500 replicates of $K=23$ quantiles were estimated from simulated data with sample sizes $n\in\{50,150,500,1{,}000,2{,}000,\\ 5{,}000\}$. QM for each replicate of simulated quantiles was done using QGP, ORD, IND, SPL, and KDE. For the QGP, ORD, and IND models, $\theta=\{\theta_{1},\theta_{2},\theta_{3},\theta_{4}\}$ where $\theta_{c}=(\mu_{c},\sigma_{c})$ for $c\in\{1,2,3,4\}$. The prior distributions assigned were $\mu_{c}\overset{ind}{\sim}N(5,7^{2})$, $\sigma_{c}\overset{ind}{\sim}N(0,6^{2})\mathds{1}\{\sigma_{c}>0\}$, $w\sim Dir(\boldsymbol{\alpha}),\quad\alpha_{c}=1$, and $\nu\sim N(0,3000^{2})\mathds{1}\{\nu>0\}$ where $\nu=n$ for the QGP and ORD models and $\nu=1/\sigma_{\rho}$ for the IND model. Often when implementing posterior sampling for a mixture distribution model, one deals with an issue called the label-switching problem. This is where for a mixture distribution parameter $\theta=\{\theta_{1},...,\theta_{C}\}$, the model likelihood is the same for different permutations of $\theta$. This lack of identifiability for elements of $\theta$ makes parameter inference useless, but the predictive distribution may still be close to the true distribution Stephens (2000). Because of this, the HMC posterior sampling chain was run longer than in the studies in the previous section. The chain was run for 80,000 steps where the first 20,000 draws were discarded as a burn-in. For this study, we were concerned only with QM and not with parameter inference, so the fact that the parameters are unidentifiable was not considered critical. However, when assessing the TV and KLD between the QM and true distributions, variational Bayes (VB) in Stan was used instead of HMC for fitting the models Kucukelbir et al. (2015). This decision was made because of the lack of parameter identifiability and because the marginal parameter means are used to estimate the PDF. While MCMC methods perform better on parameter inference, VB methods have been shown to have predictive performance comparable to MCMC Blei et al. (2017).
@@
Figures 11, 12, and 13 in the supplementary material show examples of fits of $K=23$ quantiles simulated from the three distributions for different sample sizes $n$. Included QM methods are KDE, SPL, IND, and QGP. ORD was excluded to make visualization easier, but the fits are very similar to the QGP fits. The KDE and SPL fits provide no uncertainty estimation of the quantiles, and return only a continuous function. The SPL fits show a lot of wiggle with the wiggle decreasing as $n$ increases, whereas the KDE fits do not show as much wiggle, but the KDE fits the distribution tails poorly even as $n$ increases. The IND and QGP models provide uncertainty in the fits, but the QGP is much more conservative with wider intervals which provide superior coverage of the quantiles and the PDFs.
@@
The top of figure 3 shows the 95% simulated coverage of the true quantiles for QGP, ORD, and IND fits of quantiles from the EV, La, and MIX distributions. The percent coverage is averaged over all $K=23$ quantiles and all 500 simulated replicates. The plots show that the QGP and ORD models are largely able to meet and exceed the nominal coverage where the IND model falls very short. Because it took a long time to evaluate the QF when calculating coverage, calculation was done on a thinned sample of the posterior distribution where only every $100^{th}$ HMC draw was kept so that the coverage was calculated using 6,000 posterior draws rather than the full 60,000.
@@
The bottom of figure 3 shows the UWD1, TV, and KLD distances between QM and true distributions averaged over 500 simulation replicates for the five QM methods. In the UWD1 plot, there appears to be little or no difference in performance between the QGP, ORD, IND, and SPL fits, but each of these outperforms the KDE fits. For TV, the QGP, ORD, and SPL perform similarly, though the QGP does the best for the MIX distribution. And for KLD, the QGP and ORD perform the best for lower values of $n$, but as $n$ increases, SPL appears to perform the best. Additional plots and results from these analyses are found in the supplementary material.
This section shows the ability of the QGP model to accurately perform parameter inference and QM when given sample quantiles. Although the inference for QGP is asymptotic, the QGP performance in inference and in QM is similar to that of ORD where the inference is exact. There are also circumstances where the QGP may outperform the ORD as shown in cases where the QF is easily evaluated but where the CDF is difficult to evaluate. Another case where the QGP outperformed the ORD was in QM of the normal mixture distribution. An advantage that the QGP has over SPL and KDE is that it provides accurate uncertainty estimation of quantile uncertainty where SPL and KDE provide no uncertainty estimation.

### v2

@@ 3 Quantile Gaussian process model
The covariance function $\kappa_{\theta}(\cdot,\cdot)$ initially appears as if it will be difficult to evaluate and indeed can be, depending on the functional forms of $f_{\theta}$ and $Q_{\theta}$. But where these functions belong to distributions in a location-scale family, the covariance function can be greatly simplified. Noting that the covariance is a function of $f_{\theta}(Q_{\theta}(p))=q_{\theta}(p)$, the QDF, if $\theta=(\mu,\sigma)$ where $\mu$ is a location parameter and $\sigma$ is a scale parameter, then $q_{\theta}(p)=\sigma q(p)$ where $q(p)$ is the QDF of a standard distribution having location 0 and shape 1. [49] calls this the location invariant and scale equivariant property. For location-scale distribution families, this greatly simplifies the form of $\kappa_{\theta}(\cdot,\cdot)$ so that it becomes
@@ 3.1 Normal QGP
$\Psi_{ij}=\frac{2\pi(p_{i}\wedge p_{j}-p_{i}p_{j})}{\text{exp}\{-\frac{1}{2}[\Phi^{-1}(p_{i})^{2}+\Phi^{-1}(p_{j})^{2}]\}}.$
@@
In this model, one simplifying assumption may be to treat $\sigma$ and $\gamma^{2}$ separately and proceed to fit a standard linear regression model. All frequentist and Bayesian results of the linear regression model apply, including the existence of conditionally conjugate prior distributions. The positive constraint on $\sigma$ can also be dealt with without adding much complication (see [50] pgs. 377-378). If, for a certain problem it is important to make inference on the unknown $n$, or $\sigma$ and $\gamma^{2}$ cannot treated as separate, then one may estimate the two parameters by assigning appropriate prior distributions to $\sigma$ and $n$ and reverting back to (10). In section 4, we analyze a normal QGP in a simulation study where we assign independent prior distributions to $\sigma$ and $n$, and $\sigma$ is treated as both a regression coefficient and as part of the variance.
@@ 3.2 Finite normal mixture QGP
In section 4, we analyze PIT transformed QGP model (12) where $F_{\theta}$ is set as a normal mixture distribution. Fitting this model proved faster and produced better results than trying to fit the QGP model (9) and evaluating $Q_{\theta}$ the QF of a normal mixture. In fact we found that in cases where $F_{\theta}$ was a simpler function, such as a normal or exponential distribution, fitting model (12) was at least as fast and the results were at least as good as fitting (9). Thus when possible, we elected to fit (12).
@@ 3.3 Competing quantile matching methods
The SPL method was used by [21] and [51] in order to estimate a CDF function given quantile forecasts from disease outbreak forecast hubs. SPL is an interpolation where monotonic cubic splines are fit to pass through each given quantile and predict a function for all values between given quantiles and beyond the extreme values. An R package used for fitting SPL and which we use in section 4 is distfromq [52].
The KDE method treats given quantiles as if they constitute a random sample from a distribution and applies kernel smoothing to estimate a density function. KDE requires selecting a kernel function, often a Gaussian kernel is chosen, and applying that function to each draw of a sample. [22] applied KDE smoothing with an Epanechnikov kernel to quantiles estimated via three different machine learning methods used for predicting crop yield. They then combined the estimated densities into an ensemble prediction. [26] also use KDE to estimate density forecasts from quantiles to produce energy forecasts. We use the evmix package for performing QM using KDE [53].
One of the more common methods for QM is to select a CDF or QF function and then select model parameters which give a least squares fit to the estimated quantiles as done by [23]. The R package rriskDistributions performs this least squares fit for some standard distributions [29]. Independent random error may be included to the best fit QF allowing for estimation of the variability of the quantiles [9]. In section 4, we analyze model (13) as a proxy for the IND model to compare with other methods.
@@
The final method we include is the ORD model introduced by [9]. This model relies on the definition of sample quantiles being order statistics. In the ORD model, the joint distribution of a set of order statistics is the model likelihood. The likelihood of a set of order statistics from a continuous distribution is a function of both the CDF and PDF of a continuous distribution. ORD is the method most similar to QGP and it provides exact inference of quantile uncertainty whereas the QGP provides asymptotic inference of uncertainty.
@@ 3.4 Distance measures
A number of metrics for measuring the distance between two continuous univariate distributions exist. Those we consider in this manuscript are the Wasserstein distance (WD) and a slight modification of it, the total variation (TV) or the statistical difference, and the Kullback-Leibler divergence (KLD). A brief overview of these and other metrics, some of their statistical uses and properties, and relationships between metrics is found in [54]. These metrics were used in section 4 to assess how well different QM methods approximate a true distribution.
The WD is used to measure the distance between two univariate random variables by their CDFs or QFs and has many applications in mathematics, optimization, and statistics, see [55] for a review of the WD and its use in statistics. The $p$-WD is defined for two continuous random variables $X$ and $Y$ with respective CDFs $F_{X}$ and $F_{Y}$. In terms of the two CDFs the $p$-WD is defined in (14), and the $p$-WD in terms of the corresponding QFs is defined in (15).
$WD_{p}(F_{X},F_{Y})=\left(\int_{\mathbb{R}}|F_{X}(t)-F_{Y}(t)|^{p}dt\right)^{1/p}$ (14)
$WD_{p}(F_{X},F_{Y})=\left(\int_{0}^{1}|F^{-1}_{X}(t)-F^{-1}_{Y}(t)|^{p}dt\right)^{1/p}$ (15)
We define a modified version of the WD, the uniform 1-WD (UWD1) in (16). Here we take $F_{X}$ to be the CDF of a continuous random variable $X$. For a different continuous random variable $\xi$ with CDF $F_{\xi}$, $F_{X}(\xi)$ is a random variable with support on $[0,1]$. We let $F_{X,\xi}$ be the CDF of the random variable $F_{X}(\xi)$. Then the measure in (16) measures how close $F_{X}(\xi)$ is the CDF of a standard uniform distribution with CDF $\mathcal{U}$, noting that by the PIT, if $\xi\overset{D}{=}X$, $F_{X}(\xi)$ is the random variable of a standard uniform distribution. The integral in (16) takes values between 0 and 1/2 [56], and multiplying by two ensures the measure takes values between 0 and 1. This makes the UWD1 more interpretable than the WD where a value near 0 suggests $\xi$ and $X$ are “near” each other, and a value near 1 suggests they are far apart.
$UWD_{1}(F_{X,\xi},\mathcal{U})=2\int_{0}^{1}|F_{X,\xi}(u)-x|du$ (16)
The TV, defined in (17), is a distance between distributions measured in terms of the PDFs $f$ and $g$. The TV takes values between 0 and 1, and it is closely related to the distance used in [27] to measure the goodness of QM.
$TV(f,g)=\frac{1}{2}\int_{-\infty}^{\infty}|f(x)-g(x)|dx$ (17)
@@ 4 Quantile matching methods comparison on simulated data
In this section we present simulation studies assessing the QM of the QGP model and compare results with those of the SPL, KDE, IND, and ORD matching methods. The QGP, IND, and ORD models were each fit via Hamiltonian Monte Carlo (HMC) sampling using the Stan software and the cmdstanr package developed and maintained by the [57] [58]. Simulation studies are done to assess parameter estimation and inference as well as QM estimation of a distribution.
4.1 Parameter estimation and quantile matching for known distribution families
@@ 4.1.1 Normal parameter estimation
In this simulation study, we fit the ORD and IND models to quantiles estimated from independent samples from a normal distribution and compare parameter estimation with fits from the QGP model. Quantiles were simulated by first drawing a sample of size $n$ from a known distribution then estimating $K$ quantiles given probabilities $\{p_{1},...,p_{K}\}$. The ORD, IND, and QGP models were then fit to the quantiles where $F$ was set to be the CDF of a normal distribution. For each of $n\in\{50,150,500,1{,}000,5{,}000\}$ and $K=23$, there were 500 simulation replicates. The data were simulated from a normal distribution with mean $\mu=4$ and standard deviation $\sigma=3.5$. For each of QGP, ORD, and IND models, we assigned the same prior distributions to $\mu$ and $\sigma$. When $n$ is unknown, we also assigned it a prior distribution. As one will generally have access to the quantiles being fit, it is reasonable that an informative prior, centered near where the data is centered, be assigned to $\mu$. We took this approach herein, and for scale parameters, we assigned normal priors truncated at 0 as recommended by [59]. The prior distributions were $\mu\sim N(5,7^{2})$, $\sigma\sim N(0,6^{2})\mathds{1}\{\sigma>0\}$, and $n\sim N(0,3000^{2})\mathds{1}\{n>0\}$. For the IND model the prior on the independent parameter $\sigma_{\rho}$ from model (13) is $1/\sigma_{\rho}\sim N(0,3000^{2})\mathds{1}\{\sigma_{\rho}>0\}$, which is similar to the prior on $n$ for the QGP and ORD models. For each fit, the HMC chain was run for 60,000 draws with the first 10,000 discarded as a burn-in.
The top of figure 2 shows the 90% coverage of the posterior distribution credible intervals as $n$ increases for $\mu$, $\sigma$ and $n$ for five models. QGP-n and ORD-n models where $n$ is known are included along with QGP, ORD, and IND. The 90% credible intervals of the parameters were calculated by computing the $0.05^{th}$ and $0.95^{th}$ quantiles from the posterior distribution samples. The coverage percentage is calculated as the percentage of the 500 replicates for which the true parameter value was within the 90% credible interval. As expected, the nominal coverage for the QGP and ORD models is better than that of the IND model, and for the two models where $n$ is known the coverage is better than where it is unknown, particularly for lower values of $n$ and $K$. The coverage for the ORD models appears to be slightly better than for the QGP models though not by much, and the difference decreases as $n$ increases. The superior performance of the ORD model is possibly because it provides exact inference whereas the QGP model provides asymptotic inference.
@@
To analyze how well QM methods produced predictive distributions which approximate the true distribution $N(4,3.5^{2})$, we measured the distance between the estimated predictive distributions of QGP, ORD, and IND fits to the true distribution. We also measured the distances of the SPL and KDE QM fits to the true distribution. The distances were measured via the UWD1, TV, and KLD metrics. The UWD1 for each of QGP, ORD, QGP-n, ORD-n, and IND were calculated as follows. For $\{\theta_{m}\}^{M}$ being $M$ draws from the posterior distribution of $\theta$ where $\theta_{m}=(\mu_{m},\sigma_{m})$, the QM posterior predictive distribution was simulated by repeatedly sampling a $\theta_{m}^{*}$ from $\{\theta_{m}\}^{M}$, then sampling $X_{m}$ from a distribution with CDF $F_{\theta_{m}^{*}}$. Repeating this for $M=50,000$ times gives a QM posterior predictive sample. For the CDF $F_{\theta}$ where $\theta$ is the true parameter, $\xi_{m}=F_{\theta}(X_{m})$ was calculated and the empirical CDF $\hat{F}_{\xi}$ was calculated from the sample $\{\xi_{m}\}^{M}$. $F_{\xi}$ in (16) was then replaced by $\hat{F}_{\xi}$ to calculate UWD1. For SPL and KDE $F_{\xi}(x)$ was replaced by the respective estimated CDFs. The integrate function in R was used for calculating the UWD1 in (16).
@@
where $S$ is the total number of Monte Carlo samples and $Y_{s}\overset{iid}{\sim}F_{\theta}$. For both SPL and KDE, $f_{\bar{\theta}_{M}}$ is the estimated PDF.
@@
This study shows the ability of the QGP QM model to estimate and perform inference on model parameters where the true distribution family is known. It also shows the QGP’s ability to produce a posterior predictive distribution which closely matches a true distribution, according to several metrics, relative to other QM methods. Between parameter inference and predicting the true distribution, QGP is superior to all methods we compared it with except for the ORD model which has similar results to –or for small sample size slighty better than– the QGP. However, we feel the QGP’s linear model form is easier to understand than the ORD model and thus preferable for interpretation. Additional plots for this study and for a similar simulation study where the data was simulated from an exponential family distribution are in the supplementary materials. The results of the exponential family study are similar to those of the normal distribution family study.
@@
The analysis above show the usefulness in using the QGP model both in parameter estimation and in QM when given sample quantiles. In the normal case, the QGP showed improved parameter inference over the IND model and allows for making inference on the sample size $n$. Compared to the SPL and KDE QM methods, the QGP fits tend to be closer in distance to the true distribution. The QGP model, however, performs about the same or for smaller $n$ slightly poorer than the ORD model in terms of inference and QM. However, we feel the QGP is a simpler model to understand than the ORD and easier to interpret. The following section outlines a situation where the QGP model may be a better option for QM than the ORD model.
@@ 4.1.2 Quantile defined distributions
The GLD is reviewed in section 2.1.4, and a simulation carried out similarly to the previous studies was performed with GLD as the true distribution. However, QM was only done by the QGP and ORD models. Because of the relative difficulty of evaluating the CDF of the GLD and the ease of evaluating the QF, the CDF transformed QGP model in (9) instead of (12) was used. For the ORD model, evaluating the CDF was required for modeling. Evaluating the CDF of the GLD requires using an algebraic solver, and it took longer to code the ORD model in Stan due to a lack of readily available software which we were aware of. Once working, however, both QGP and ORD models fit the estimated quantiles well, however, the ORD usually required more time to fit than the QGP. Figure 3 shows boxplots for the 500 replicates for different values of $K$ quantiles. In all cases except for $K=50$ the QGP tends to fit much faster than ORD. Note that the $y$-axis is on the logarithmic scale. Even at $K=50$ some outliers for the ORD model required hundreds of minutes to fit whereas the longest required time for fitting a QGP model was a few dozen minutes.
Figure 3: Boxplots of time to fit a model in seconds ($y$-axis) for the 500 replicates in the simulation study for QM of the generalized lambda distribution (GLD) for $K\in\{3,7,15,23,50\}$ ($x$-axis). Boxplots are separated by QM methods QGP (red) and ORD (blue). The $y$-axis is on the $\text{log}_{10}$ scale.
@@ 4.2 Matching unknown distributions
With the true distribution family being unknown, we made the modeling decision to set $F$ in the QGP model of (12) to be a mixture of normals distribution taking the form of (11). This decision to use a normal mixture was based on the common claim that any continuous distribution may be approximated by a finite mixture of normal distributions [60, 61, 62]. This claim may be optimistic, but we found that with enough components, the normal mixture distribution can approximate the non-normally shaped distributions reasonably well. However, any function which meets the technical definition of a CDF would be appropriate to use for $F$, see [39] for an example. After a small analysis comparing fits of one to six normal components, the number of component distributions was chosen to be $C=4$. The results, included in the supplementary materials, show that as the number of components in the mixture distrubtion increases, the distribution approximation improves. But for the cases considered, the improvement from four to five components or more is small.
@@
The three distributions from which data were simulated and sample quantiles estimated and to which QM methods were fit were the extreme value (EV) distribution with location 0 and scale 1 $EV(0,1)$, Laplace (La) with location 0 and scale 1 $La(0,1)$, and a two component normal mixture (MIX) wN(1,0.9)+(1w)N(1.2,0.6)wN(-1,0.9)+(1-w)N(1.2,0.6) where $w=0.35$. From the three distributions, 500 replicates of $K=23$ quantiles were simulated for each of $n\in\{50,150,500,1{,}000,2{,}000,5{,}000\}$. QM for each replicate of simulated quantiles was done using QGP, ORD, IND, SPL, and KDE. For the QGP, ORD, and IND models, $\theta=\{\theta_{1},\theta_{2},\theta_{3},\theta_{4}\}$ where $\theta_{c}=(\mu_{c},\sigma_{c})$ for $c\in\{1,2,3,4\}$. The prior distributions assigned were $\mu_{c}\overset{ind}{\sim}N(5,7^{2})$, $\sigma_{c}\overset{ind}{\sim}N(0,6^{2})\mathds{1}\{\sigma_{c}>0\}$, $w\sim Dirichlet(\boldsymbol{1}_{C})$, and $\nu\sim N(0,3000^{2})\mathds{1}\{\nu>0\}$ where $\nu=n$ for the QGP and ORD models and $\nu=1/\sigma_{\rho}$ for the IND model. Often when implementing posterior sampling for a mixture distribution model, one deals with an issue called the label-switching problem. This is where for a mixture distribution parameter $\theta=\{\theta_{1},...,\theta_{C}\}$, the model likelihood is the same for different permutations of $\theta$. This lack of identifiability for elements of $\theta$ makes parameter inference useless, but the predictive distribution may still be close to the true distribution [63]. Because of this, the HMC posterior sampling chain was run longer than in studies of the previous section. The chain was run for 80,000 steps where the first 20,000 draws were discarded as a burn-in. For this study, we were concerned only with QM and not with parameter inference, so the fact that the parameters are unidentifiable was not considered critical. However, when assessing the TV and KLD between the QM and true distributions, variational Bayes (VB) in Stan was used instead of HMC for fitting the models [64]. This decision was made because of the lack of parameter identifiability and because the marginal parameter means are used to estimate the PDF. While MCMC methods perform better on parameter inference, VB methods have been shown to have predictive performance comparable to MCMC [65].
@@
Figure 4, shows examples of fits of $K=23$ quantiles simulated from the EV distribution for different sample sizes $n$. Included QM methods are KDE, SPL, IND, and QGP. ORD was excluded to make visualization easier, but the fits are very similar to the QGP fits. The KDE and SPL fits provide no uncertainty estimation of the quantiles, and return only a continuous function. The SPL fits show a lot of wiggle with the wiggle decreasing as $n$ increases, whereas the KDE fits do not show as much wiggle, but the KDE fits the distribution tails poorly even as $n$ increases. The IND and QGP models provide uncertainty in the fits, but the QGP is much more conservative with wider intervals which provide superior coverage of the quantiles and the PDFs.
@@
The top of figure 5 shows the 95% simulated coverage of the true quantiles for QGP, ORD, and IND fits of quantiles from the EV, La, and MIX distributions. The percent coverage is averaged over all $K=23$ quantiles and all 500 simulated replicates. The plots show that the QGP and ORD models are largely able to meet and exceed the nominal coverage where the IND model coverage is low. Because of the time it took to evaluate the QF when calculating coverage, calculation was done on a thinned sample of the posterior distribution where only every $100^{th}$ HMC draw was kept so that the coverage was calculated using 6,000 posterior draws rather than the full 60,000.
@@
The bottom of 5 shows the UWD1, TV, and KLD distances between QM and true distributions averaged over 500 simulation replicates for the five QM methods. In the UWD1 plot, there appears to be little or no difference in performance between the QGP, ORD, IND, and SPL fits, but each of these outperforms the KDE fits. For TV, the QGP, ORD, and SPL perform similarly, though the QGP does the best for the MIX distribution. And for KLD, the QGP and ORD perform the best for lower values of $n$, but as $n$ increases, SPL appears to perform the best. Additional plots and results from these analyses are found in the supplementary materials.
This section shows the ability of the QGP model to accurately perform parameter inference and QM when given sample quantiles. Although the inference for QGP is asymptotic, the QGP performance in inference and in QM is similar to that of ORD which provides exact inference. There are also circumstances where the QGP may outperform the ORD as shown in cases where the QF is easily evaluated but where the CDF is difficult to evaluate. Another case where the QGP outperformed the ORD was in QM of the normal mixture distribution. An advantage that the QGP has over SPL and KDE is that it provides accurate uncertainty estimation of quantile uncertainty where SPL and KDE provide no uncertainty estimation.